OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x8 + 3x6) cubic feet. What polynomial represents the height of the box in feet?

OpenStudy (anonymous):

4x6+3x4

OpenStudy (anonymous):

we have Ah = 4x8+3x6 A=x2= and the area of ​​the rectangle of the base of the box h=the height of the box

OpenStudy (anonymous):

so that would come out to be 4x6 + 3x4

OpenStudy (anonymous):

Where does the x2 come in like what do you do with that?

OpenStudy (anonymous):

The area is x • x = x^2 Area • height = volume Height = volume / area

OpenStudy (anonymous):

thank you @TURITW

OpenStudy (anonymous):

Oh I understand. Thanks guys! :D

OpenStudy (anonymous):

you're welcome!!